OpenStudy (anonymous):

MEDALS! MEDALS MEDALS!! If ∠m ≅ ∠p, m∠m = 9x - 2, and m∠p = 4x + 18, how many degrees are in ∠p?

OpenStudy (anonymous):

Given: ∠m ≅ ∠p, hence: 9x - 2 = 4x + 18 9x- 4x= 20 5x= 20 x= 4 m∠p = 4x + 18 = 4*4 +18 = 16 +18 = 34 degrees